# Avoiding matrix exponentials for large transition rate matrices

### 09 Dec 2023 v1

@@ 3.2.1 Time comparison
To compare the time needed to calculate the final distribution (thus avoiding the matrix exponential in (2)) using for each method, we measure the wall time required to sample 100 values for all kinetic parameters from across datasets (each demanding different total size of state space). To obtain some statistics, we repeat this procedure 20 times. The results can be seen in Fig. 3. We focused solely on comparing the time aspects, conducting tests under conditions where MCMC sampler had already converged to the posterior by initializing it at the simulation’s ground truth.
@@
Our Python code, available in our repository [39], leverages the Numba compilation tool [41] and a package created by one of us that enables sparse matrices in Numba-compiled code [42]. The benchmark was conducted on a system equipped with an Intel(R) Core(TM) i7-7700K CPU operating at a base frequency of 4.20 GHz. The system is outfitted with 16 GB of RAM.
@@
Figure 3: We recorded the wall time, $t_{W}$, necessary to obtain 100 samples from the posterior in relation to the total number of states. A comparison of RK4, Krylov, J-MJP, and R-MJP for all three examples is presented above. In the birth-death process (Sec. 3.1.1) is presented on the left panel, the two states birth-death (Sec. 3.1.2) is presented in the middle panel example and for the autoregulatory gene model (Sec. 3.1.3) is presented in the right panel. In the birth-death example, we also compare to the conventional implementation of the matrix exponential (ME). When considering overall performance, the contest narrows down between R-MJP and Krylov, with Krylov tends to exhibit superior performance, especially as the state spaces increase in size.
@@
Fig. 3 clearly illustrates the expected quadratic scaling of RK4, J-MJP, and R-MJP relative to the total number of states for the birth-death and two states birth-death. In these cases, we observe a slightly smaller exponent due to necessary overhead. Nevertheless, this overhead is anticipated to diminish in significance with an increase in state space size. The Krylov scaling is somewhat less straightforward as the method has a more complex dependence on the number of states, as discussed in Sec. 2.2.5 and Table 1, as such needs a considerably larger number of states to actually reach a regime of quadratic scaling. When studying the autoregulatory gene system, the exponents of the power law relating the wall time and number of states are smaller due to the less straightforward relationship between the largest element of $\mathbf{A}_{\theta}$ and the total number of states, as mentioned in Sec 3.1.3. In the same figure we also show, for the birth-death process Sec. 3.1.1, how all of these methods dramatically outperform regular matrix exponential implementation. This was not implemented in the two other systems, as the difference is even more dramatic. The competition for overall best method becomes between R-MJP and Krylov, with the tendency of Krylov to better perform at larger state spaces.
@@ 3.2.2 J-MJP is inefficient
As mentioned in Sec. 2.1.2 the samplers generated by #3, #4, and #5 obtain samples from the full posterior $p(\theta|\bar{s})$. Consequently, the samplers from Methods #3, #4, and #5 recover nearly identical probability distributions over model parameters $\theta$, with the only variation arising from the approximation error in the likelihood.
Meanwhile, Method #2 uses the set of number of virtual jumps $\bar{k}$ as a latent variable. To effectively compare the performance of J-MJP (Method #2) to other methods, it is essential to look not only at execution time but also to assess how well an equivalent number of samples describe the posterior. This become specially important since sampling $\bar{k}$ introduces extra variables equal to the number of observations. This inclusion dramatically increases the dimensionality of the sample space. With this increased dimensionality, it’s logical to assume that more samples may be required to properly characterize the joint posterior between $\theta$ and $\bar{k}$.
In Fig. 5, the posteriors obtained by J-MJP and the other three benchmarked methods are shown to give inconsistent results in both learned parameters and their uncertainty (depicted as confidence intervals in 5) . This comparison was done for the two states birth-death example (Sec. 3.1.2) with ground truth kinetic parameters $\theta=\{\beta,\gamma,\lambda_{\text{act}},\lambda_{\text{ina}}\}=\{200,1,2,1\}$. More examples in Supplemental Information B.
Figure 4: Histogram of the posterior samples of each kinetic parameter for the autoregulatory gene networks. While R-MJP, RK4, and Krylov yield coherent posteriors, J-MJP, while placing the bulk of its probability density in a similar parameter region, recovers confidence intervals inconsistent with the other methods.
Figure 4: Histogram of the posterior samples of each kinetic parameter for the autoregulatory gene networks. While R-MJP, RK4, and Krylov yield coherent posteriors, J-MJP, while placing the bulk of its probability density in a similar parameter region, recovers confidence intervals inconsistent with the other methods.
Figure 5: Auto-correlation function (30) for the kinetic parameters sampled from J-MJP and R-MJP within the autoregulatory gene network. It is important to notice that this confirms that the sampler generated using J-MJP is inefficient, as the autocorrelation does not move to near 0 even after tens of thousands samples, while in R-MJP (similar behaviour is observed for Krylov and RK4) the autocorrelation falls to zero after just a few hundred samples.
@@
with $\theta^{a}_{h}$ being the $a$-th parameter of the $h$-th sample and $\bar{\theta}^{a}=\sum\limits_{h=1}^{H}\theta^{a}_{h}$. In simple terms, $\text{ACF}(\tau)$ quantifies the average correlation between a sample and another with another sample $\tau$ steps away within the Markov chain. Here it is important to mention that the MCMC structure is intrinsic to J-MJP, while the marginalized methods can be used to calculate the vector times matrix exponential (2), both within an MCMC framework and independently.
An ideal scenario would have ACF rapidly descending from one to zero, signifying optimal mixing. Nonetheless, Fig. 5 reveals that while this is achieved by R-MJP (and the other marginalized methods) for J-MJP sampled kinetic parameters remain correlated, even following thousands of iterations. Consequently, at least an order of magnitude more samples are needed to characterize the posterior as proficiently with J-MJP as with alternative methods. This happens as a consequence of J-MJP having to sample a considerable number of latent variables. Thus a method only sampling kinetic parameters will, consequentially, more exhaustively sample its posterior in fewer iterations.
4 Discussion
Our comparison of different methods to avoid matrix exponentiation reveals that no one method is always preferable. Fig. 3 indicates that the most efficient method to generate sufficient samples of the posterior hinges on both the system and the state space: either the R-MJP (Method #3) or the Krylov subspace (Method #5), with the latter typically proving more efficient for significantly larger state spaces and the former seeming to have advantage for sparser matrices. However, the other methods should not be dismissed entirely for reasons we elaborate now.
Although RK4 (Method #4) performed significantly worse than all other methods except the complete matrix exponential, it was the only method generalizable to cases where transition rates change in time – such as induction [43] – since it doesn’t invoke a time-constant generator matrix.
@@ 4 Discussion
Opting for synthetic data was crucial to demonstrate, in a controlled environment, the wall time scaling across state spaces size ranging in the hundreds to thousands of states. Our findings suggest that, for practical purposes, the choice between Krylov and R-MJP is highly contingent, thus recommending a benchmark tailored to the specific dynamics under study.
Furthermore, although the time benchmark indicates that Krylov outperforms R-MJP in large state spaces, it is pivotal to note that it demands more memory allocation, associated with the choice of subspace size, $\kappa$. In contrast, R-MJP does not require any such choices. The effective time step, $1/\Omega_{\theta}$, is determined by the system, and the selection of $k^{\ast}$ establishes a directly computed (and optimistically presumed) upper bound on the error.
@@ Acknowledgments
Supplemental Information
@@ Appendix A Sampling scheme for the models in examples
In this Supplemental Information section, we describe how to sample the posterior the Bayesian posterior, as mentioned in Sec. 2.1.2. First we will give an overview of the specific strategy for the MCMC based sampler known as Metropolis-Hastings (MH) and how it was applied to the examples mentioned in the main text. In Sec A.2 we describe the form of adaptive MH used. In Sec A.3 we expand upon the extra work necessary to sample $\bar{k}$ in method #2. Finally, in sec A.4 we give some more detail for the prior, Monte Carlo chain initialization scheme, and the parameters used in the adaptive scheme.
@@ A.1 Metropolis-Hastings sampling scheme
In order to create a MH based Markov chain Monte Carlo sampler that obtains $\theta$ from the posterior $\theta\sim p(\theta|\bar{s})$ — although the method described here can be used for sampling from any distribution — one has to initialize a sequence with a value $\theta_{0}$ and obtain a sequence (or chain) of samples $\{\theta_{0},\theta_{1},\ldots\theta_{H}\}$. With each iteration, from $\theta_{h}$ to $\theta_{h+1}$ following the steps:
@@
This ensures that the sample chain converges to the desired target distribution $p(\theta|\bar{s})$ [3]. In principle it is independent of the proposal distribution $q$ and the initialization $\theta_{0}$, nevertheless these should be carefully selected in order to have an efficient sampler — meaning a sampler converging fast and exploring a significant area of the target distribution with fewer samples.
A usual form of the proposal distribution when dealing with dynamical processes is to treat the set of kinetic parameters as a vector $\theta\in\mathbb{R}^{\eta}$ with $\theta=(\theta^{1},\theta^{2},\ldots,\theta^{\eta})$ and proposing from a multivariate normal distribution in log space
@@
where $\Psi_{h}=(\Psi^{a}_{h})=(\log\theta_{h}^{a})$. With $\Psi_{\text{prop}}$ sampled from (32), we obtain $\theta^{a}_{\text{prop}}=e^{\Psi^{a}}$. The covariance matrix $\mathbf{Z}$ is initially set as $\mathbf{Z}=z\mathbf{I}$, where $z$ is a small positive value and $\mathbf{I}$ is the $\eta\times\eta$ identity matrix. For an efficient sampler for the data at hand, it is useful to find adapt $\mathbf{Z}$, meaning changing $\mathbf{Z}$ to values that generate an efficient distribution., we explain how this is done in the following subsection.
@@ A.2 Adaptive Metropolis-Hastings
    $\begin{split}\mathbf{Z}&=\left(\frac{2.38}{\eta}\right)^{2}\left[\text{cov}({\theta_{H-H^{\prime}},\theta_{H-H^{\prime}+1},\ldots,\theta_{H}})+\varepsilon\mathbf{I}\right]\ ,\\ \text{where}&\ \text{cov}({\theta_{1},\theta_{2},\ldots,\theta_{H^{\prime}}})=\left(\frac{1}{H^{\prime}}\sum_{h=1}^{H^{\prime}}\theta^{T}_{h}\theta_{h}\right)-\bar{\theta}^{T}\bar{\theta}\ ,\end{split}$ (33)
@@
This procedure is repeated several times, after such the MH sample chain restarts using the final $\mathbf{Z}$ obtained from the adaptation process, effectively only the samples obtained after stabilizing $\mathbf{Z}$ are saved as samples from the posterior. The choice of all parameters is explained in Sec. A.4.3.
@@ A.3 Sampling jumps in J-MJP
When sampling from method #4 in (2.2.2) we also apply a MH strategy. When sampling $\theta$ from $p(\theta|\bar{s},\bar{k})$ we use the sampling scheme outlined in Sec. A.1. When sampling $\bar{k}$ from $p(\bar{k}|\bar{s},\theta)$we use a propose each $k^{i}$ separately as
$q(k^{i}_{\text{prop}}|k^{i}_{h})=P_{\text{Poisson}}(k^{i}_{\text{prop}}|\Omega_{\theta}T^{m})\ .$ (34)
@@ A.4.1 Prior
To compute the Bayesian posterior $p(\theta|\bar{s})$ as given by Equation (11), we require two components: the prior $p(\theta)$ and the likelihood $p(\bar{s}|\theta)$. While the discussion in Sec. 2.2 presents the likelihood, the choice of prior is tailored to the specific problem under consideration. Since our objective here was to characterize the different methods, we opt to use a prior as uniform as possible while keeping an analytic function for safe implementation within the multiple algorithms. Thus for all inference processes studied here we choose the prior
$p(\ln\theta)=\prod_{a=1}^{\eta}\frac{1}{\sigma\sqrt{2\pi}}\exp\left(-\frac{(\ln\theta^{a}-\mu)^{2}}{2\sigma^{2}}\right)\ ,$ (36)
with $\mu=1$ and $\sigma=10^{5}$ Given that $\sigma$ is significantly large, the prior exerts only a minimal influence on each individual kinetic parameter. Specifically, for $\theta^{a}$ values ranging from $0.01$, the smallest ground-truth kinetic parameter in our study, to $800$, the largest ground-truth value, the prior remains uniform up to four decimal places.

### 06 Feb 2024 v3

@@ 3.2.1 Time scaling comparison
To compare the time needed to calculate the final distribution (thus avoiding the matrix exponential in (2)) used for each method, we measure the wall time required to sample 100 values for all kinetic parameters from across datasets (each demanding different total state space size). To obtain some statistics, we repeat this procedure 20 times. The results can be seen in Fig. 3. We focused on comparing the time scaling between methods. conducting tests under conditions where MCMC samplers had already converged to the posterior by initializing it at the simulation’s ground truth.
@@
Our Python code, available in our repository [36], leverages the Numba compilation tool [42] and a package enabling sparse matrix manipulation in Numba-compiled code [35]. The benchmark was conducted on a system equipped with an Intel(R) Core(TM) i7-7700K CPU operating at a base frequency of 4.20 GHz. The system is outfitted with 16 GB of RAM.
@@
Figure 3: We recorded the wall time, $t_{W}$, necessary to obtain 100 samples from the posterior in relation to the total number of states. A comparison of J-MJP, R-MJP, RK4, and Krylov (KRY) for all three examples is presented above. In the birth-death process (Sec. 3.1.1) is presented on the top panel, the two states birth-death (Sec. 3.1.2) is presented in the middle panel example and for the autoregulatory gene model (Sec. 3.1.3) is presented in the bottom panel. In the birth-death example, we also compare to the conventional implementation of the matrix exponential (ME). When considering overall performance, the contest narrows down between R-MJP and Krylov, with Krylov tends to exhibit superior performance, especially as the state spaces increase in size.
@@
Fig. 3 clearly illustrates the expected quadratic scaling of RK4, J-MJP, and R-MJP relative to the total number of states in each of the three examples under study. In these cases, we observe a slightly smaller exponent due to necessary overhead. Nevertheless, this overhead is anticipated to diminish in significance with an increase in state space size. The Krylov scaling is somewhat less straightforward as the method has a more complex dependence on the number of states, as discussed in Sec. 2.2.5 and Table 1, as such needs a considerably larger number of states to actually reach a regime of quadratic scaling.
When studying the autoregulatory gene system, the exponents of the power law relating the wall time and number of states are smaller due to the less straightforward relationship between the largest element of $\mathbf{A}_{\theta}$ and the total number of states, as mentioned in Sec 3.1.3. In the same figure we also show, for the birth-death process Sec. 3.1.1, how all of these methods dramatically outperform regular matrix exponential implementation.
The comparison between the time scaling between Methods #2 – #5 and the regular matrix exponential was not implemented in the two other systems, as the difference is even more dramatic. The competition for overall best method comes between R-MJP and Krylov, with the tendency of Krylov to better perform at larger state spaces. This difference in performance can be attributed to the fact that, in smaller state spaces, the time required for allocating memory to store the Krylov subspace (as detailed in Algorithm 3) and for exponentiating the $\kappa\times\kappa$ projection matrix, $\mathbf{H}$, is significant compared to the time needed for the vector-matrix product computation. However, as the state space expands, the relative time consumed for memory allocation and matrix exponentiation diminishes, making the Krylov method increasingly more efficient. Nevertheless, despite the size of the state space, Krylov still requires arbitrary choices of the subspace size, $\kappa$ and the time step (with the latter being also required in RK4). In contrast, R-MJP (as well as T-MJP and J-MJP) have an the effective time step, $1/\Omega_{\theta}$, endowed by the system.
Determining the state space size at which Krylov becomes more efficient than R-MJP depends on the matrix’s sparcity structure and the largest element of the matrix $\max\limits_{n}{|a_{nn}|}$ (related to the effective time step in each of the methods). Moreover, in Sec. 3.2, we focused on the time scaling of each method which, when implemented correctly, is independent of the hardware. However for the examples under investigation, using our code[36] and hardware (described above) the critical size at which Krylov becomes more efficient is when the R-MJP and Krylov lines in Fig. 3 intercept.
@@ 3.2.2 J-MJP is inefficient
As mentioned in Sec. 2.1.2 the samplers generated by #3, #4, and #5 obtain samples from the full posterior $p(\theta|\bar{s})$. Consequently, the samplers from Methods #3, #4, and #5 recover nearly identical probability distributions over model parameters $\theta$.
Meanwhile, Method #2 uses the set of number of virtual jumps $\bar{k}$ as a latent variable. To effectively compare the performance of J-MJP (Method #2) to other methods, it is essential to look not only at execution time but also to assess how well an equivalent number of samples describes the posterior. This becomes especially important since sampling $\bar{k}$ introduces extra variables equal to the number of observations. This inclusion dramatically increases the dimensionality of the sample space. With this increased dimensionality, more samples may be required to properly characterize the joint posterior between $\theta$ and $\bar{k}$.
In Fig. 4, the posteriors obtained by J-MJP and the other three benchmarked methods are shown to give inconsistent results in both learned parameters and their uncertainty (depicted as credible intervals in 4). This comparison was done for the two state birth-death example (Sec. 3.1.2) with ground truth kinetic parameters $\theta=\{\beta,\gamma,\lambda_{\text{act}},\lambda_{\text{ina}}\}=\{200,1,2,1\}$. More examples in the Supplemental Information B.
Figure 4: Histogram of the posterior samples of each kinetic parameter for the two state birth-death. The histogram samples are compared with the simulation’s ground truth (parameter values used for generating synthetic data). In each of these histograms we highlight the estimated expected value of the posterior (corresponding to the sample mean of each parameter) as well as the estimated $50\%$ and $95\%$ credible intervals (corresponding to the samples’ 25th to 75th percentiles and 2.5th to 97.5th percentiles, respectively). We notice that R-MJP, RK4, and Krylov yield coherent posteriors. On the other hand J-MJP, while placing the bulk of its probability density in a similar parameter region, recovers confidence intervals inconsistent with the other methods.
Figure 5: Auto-correlation function (30) for the kinetic parameters sampled using J-MJP (on top) and R-MJP (in the bottom) within the two state birth-death example. The results confirm that the sampler generated using J-MJP is inefficient, as the autocorrelation does not move to near 0 even after tens of thousands samples, while in R-MJP the autocorrelation falls to zero after just a few hundred samples (as occurs with Krylov and RK4).
@@
with $\theta^{a}_{h}$ being the $a$-th parameter of the $h$-th sample and $\bar{\theta}^{a}=\sum\limits_{h=1}^{H}\theta^{a}_{h}$. In simpler terms, $\text{ACF}(\tau)$ quantifies the average correlation between a sample and another with another sample $\tau$ steps away within the Markov chain. Indeed, the MCMC structure is intrinsic to J-MJP (and T-MJP) while the other three methods (R-MJP, RK4, and Krylov) calculate the vector times matrix exponential (2), both within an MCMC framework and independently.
An ideal scenario would have ACF rapidly dropping from one to zero, signifying optimal mixing. Nonetheless, Fig. 5 reveals that while this is achieved by R-MJP (as well as RK4 and Krylov), for J-MJP sampled kinetic parameters remain correlated, even following thousands of iterations. Consequently, at least an order of magnitude more samples are needed to characterize the posterior as proficiently with J-MJP as with RK4, Krylov and . This happens as a consequence of J-MJP having to sample a considerable number of latent variables.
4 Conclusion
Our comparison of different methods to avoid matrix exponentiation reveals that no one method is always preferable. Fig. 3 indicates that the most efficient method to generate sufficient samples of the posterior hinges on both the system and the state space: either the R-MJP (Method #3) or the Krylov subspace (Method #5), with the latter typically proving more efficient for significantly larger state spaces and the former seeming to have advantage for sparser matrices. However, the other methods should not be dismissed entirely for reasons we discuss below. Fig. 6 summarizes the discussion by giving a recommendation for when to use each method.
Figure 6: Summary of recommendations for when to use each of method described in the present article based on the results reported in Sec. 3.2.
Although RK4 (Method #4) performed significantly worse than all other methods, except the complete matrix exponential, it was the only method generalizable to cases where transition rates change in time – such as induction [43] – since it doesn’t invoke a time-constant generator matrix.
@@ 4 Conclusion
Opting for synthetic data was crucial to demonstrate, in a controlled environment, the wall time scaling across state spaces size ranging in the hundreds to thousands of states. Our findings suggest that, for practical purposes, the choice between Krylov and R-MJP is highly contingent. As such, we recommend a benchmark of these two methods tailored to the specific dynamics under study for anyone who needs their algorithm to run as efficiently as possible.
Furthermore, although the time benchmark indicates that Krylov outperforms R-MJP in large state spaces, Krylov still demands more memory allocation, associated with the choice of subspace size, $\kappa$. In contrast, and this is absolutely critical to emphasize, R-MJP does not require any such choices. The effective time step, $1/\Omega_{\theta}$, is determined by the system, and the selection of $k^{\ast}$ establishes a directly computed (and optimistically presumed) upper bound on the error.
@@ Acknowledgments
Supplemental Information to “Avoiding matrix exponentials for large transition rate matrices”
[…]
${}^{1}$ Center for Biological Physics, Arizona State University
${}^{2}$ Department of Physics, Arizona State University
${}^{3}$ School of Molecular Sciences, Arizona State University
@@ Appendix A Sampling scheme for the models in examples
In this Supplemental Information section, we describe how to sample the posterior the Bayesian posterior, as mentioned in Sec. 2.1.2. First we will give an overview of the specific strategy for the MCMC based sampler known as Metropolis-Hastings (MH) and how it was applied to the examples mentioned in the main text. In Sec A.2 we describe the form of adaptive MH used. In Sec A.3 we expand upon the extra work necessary to sample $\bar{k}$ in method #2. Finally, in Sec. A.4 we give some more detail for the prior, Monte Carlo chain initialization scheme, and the parameters used in the adaptive scheme.
@@ A.1 Metropolis-Hastings sampling scheme
In order to create a MH based Markov chain Monte Carlo sampler drawing $\theta$ from the posterior $\theta\sim p(\theta|\bar{s})$ as presented in Sec. 2.1.2, although the method described here can be used for sampling from any distribution. First, we ought to initialize a sequence with a value $\theta_{0}$ and obtain a sequence (or chain) of samples $\{\theta_{0},\theta_{1},\ldots\theta_{H}\}$. With each iteration, from $\theta_{h}$ to $\theta_{h+1}$ following the steps:
@@
This ensures that the sample chain converges to the desired target distribution $p(\theta|\bar{s})$ [3]. In principle it is independent of the proposal distribution $q$ and the initialization $\theta_{0}$, nevertheless these should be carefully selected in order to have an efficient sampler — meaning a quickly converging sampler exploring a significant area of the target distribution with as few samples as possible.
A usual form of the proposal distribution when dealing with dynamical processes is to treat the set of kinetic parameters as a vector $\theta=(\theta^{1},\theta^{2},\ldots,\theta^{\eta})$ and proposing from a multivariate normal distribution in log space
@@
where $\Psi_{h}=(\Psi^{a}_{h})=(\log\theta_{h}^{a})$. With $\Psi_{\text{prop}}$ sampled from (32), we obtain $\theta^{a}_{\text{prop}}=e^{\Psi^{a}}$. The covariance matrix $\mathbf{Z}$ is initially set as $\mathbf{Z}=z\mathbf{I}$, where $z$ is a small positive value and $\mathbf{I}$ is the $\eta\times\eta$ identity matrix. For an efficient sampler for the data at hand, it is useful to find adapt $\mathbf{Z}$, meaning changing $\mathbf{Z}$ to values generating an efficient distribution. We explain how this is done in the following subsection.
@@ A.2 Adaptive Metropolis-Hastings
    $\begin{split}\mathbf{Z}&=\left(\frac{2.38}{\eta}\right)^{2}\left[\text{cov}({\theta_{H-H^{\prime}},\theta_{H-H^{\prime}+1},\ldots,\theta_{H}})+\varepsilon\mathbf{I}\right]\ ,\\ \text{where}&\ \text{cov}({\theta_{1},\theta_{2},\ldots,\theta_{H^{\prime}}})=\left(\frac{1}{H^{\prime}}\sum_{h=1}^{H^{\prime}}\theta^{T}_{h}\theta_{h}\right)-\bar{\theta}^{T}\bar{\theta}\ .\end{split}$ (33)
@@
This procedure is repeated several times, more details as well as the choice of parameters in each example are found in Sec. A.4.3, following which the MH sample chain restarts using the final $\mathbf{Z}$ obtained from the adaptation process. Only the samples obtained after stabilizing $\mathbf{Z}$ are saved as samples from the posterior.
@@ A.3 Sampling jumps in J-MJP
When sampling from method #4 in (2.2.2) we also apply a MH strategy. When sampling $\theta$ from $p(\theta|\bar{s},\bar{k})$, we use the sampling scheme outlined in Sec. A.1. When sampling $\bar{k}$ from $p(\bar{k}|\bar{s},\theta)$ we use a proposal for each $k^{i}$ separately as
$q(k^{i}_{\text{prop}}|k^{i}_{h})=P_{\text{Poisson}}(k^{i}_{\text{prop}}|\Omega_{\theta}T^{m})$ (34)
@@ A.4.1 Prior
To compute the Bayesian posterior $p(\theta|\bar{s})$ as given by Equation (11), we require two components: the prior $p(\theta)$ and the likelihood $p(\bar{s}|\theta)$. While the discussion in Sec. 2.2 presents the likelihood, the choice of prior is tailored to the specific problem under consideration. Since our objective here was to characterize the different methods, we opt to use a prior as uniform as possible while keeping an analytic function for safe implementation within the multiple algorithms. Thus for all inference processes studied here we choose a normal prior for the logarithm of the kinetic parameters, $\Psi$ as defined in the paragraph containing equation (32),
$p(\Psi)=\prod_{a=1}^{\eta}\frac{1}{\sigma\sqrt{2\pi}}\exp\left(-\frac{(\psi^{a}-\mu)^{2}}{2\sigma^{2}}\right)\ .$ (36)
with $\mu=1$ and $\sigma=10^{5}$. Given that $\sigma$ is significantly large, the prior exerts only a minimal influence on each individual kinetic parameter. Specifically, for $\theta^{a}$ values ranging from $0.01$, the smallest ground-truth kinetic parameter in our study, to $800$, the largest ground-truth value, the prior remains roughly uniform up to four decimal places.